Keithf1:

If you tripled the slant height and radius of a cone, what would be the formula to find the modified surface area? A. 2004-06-01-03-00_files/i0120000.jpg B. 2004-06-01-03-00_files/i0120001.jpg C. 2004-06-01-03-00_files/i0120002.jpg D. 2004-06-01-03-00_files/i0120003.jpg

justjm:

Hi, it would be helpful if you post the pictures, too. However, I can lead you given through with the question. For any cone, the formula for the surface area can be determined by this:  Area of the base + lateral area, so \(SA=\pi r^2+\pi rs \) Let's denote \(r_1\) for radius, \(s_1\) for slant height, and \(SA_1\) for surface area all prior to modifications. Now you are tripling the radius and slant height, so your new values would be \(r_2=3r_1\) \(s_2=3s_1\) Now you can plug them into the SA formula: \(SA=πr^2+πrs\) Substitute \(SA_2=\pi (3r_1)^2+\pi (3r_1)(3s_1)\)\) Simplify \(SA_2=9\pi r_{1}^2+9\pi r_1s_1\) \(SA_2=9(\pi r_1^2+\pi r_1s_1)\) You can now substitute in the original S.A. (\(SA_1=\pi r_1^2+\pi r_1s_1 \)) \(SA_2=9(SA)\) \(SA_2=9(\pi r_1^2+\pi r_1s_1)\)